A meter stick at rest makes an angle of 30.0° with the x-axis. Find the length of the meter stick and the angle it makes with the x'-axis for an observer moving parallel to the x-axis at a speed of 0.650c
1
Expert's answer
2020-10-02T07:11:25-0400

Considering length of the stick as the Hypotenuse of Right Angled Triangle, we have-

Length of stick along x-axis = 1"\times" cos30°

= "\frac{\sqrt{3}}{2}" m

Since motion is along x-direction, length contraction will also take place along x direction only. Therefore,

lx = lox"\sqrt{1-\frac{v^2}{c^2}}"

= lox"\sqrt{1-0.4225}"

= lox (0.76)

= 0.658m

ly = loy

= sin30°

= 0.5m

Length of stick in the moving frame-

= "\sqrt{0.5^2+0.658^2}"

= 0.8264 m

The 1m long stick is contracted to 0.8264m in the moving frame

Now,

Angle stick make with x-axis-

"\theta" = tan-1("\frac{0.5}{0.658}" )

= 37.23°